Shift Summary: Day shift 7/15/1999

Tony Barker (tonyb@cuhep.Colorado.EDU)
Thu, 15 Jul 1999 14:39:34 -0600 (MDT)

Shift Summary for day shift, 07/15/99

On Shift: Ron Ray, Tony Barker

Hours of beam available at all intensities: 7:15
Hours of beam available at nominal intensity: 7:15
Hours of E832 data (special runs not included): 7:00
Hours of special runs (laser scans, muon runs...): 0:05

Accounting of lost time: change tapes; stuck busies

========================================================================
Runs ended this shift:

Run 13912 - BEAM_832 (16.4M events) to do laser scan (beam off)
Run 13913 - LASER_SCAN (aborted during first pass)
Run 13914 - BEAM_832 (8.5M events) due to full tapes.

Tapes brought to FCC this shift: none

========================================================================

Accesses during shift: none

========================================================================

Overview of shift: 832 data taking at ~6.8 E12

Smooth shift generally.
Note that fpedcal FERA pedestal monitoring histograms will not load.

========================================================================

Description of Shift:

run spill time comment
------------------------------------------------------------------------
13912 07:18 End run to do long laser scan since beam is
off.

13913 07:25 Beam returns; abort laser scan during 1st pass.

13914 001 07:40 Start new BEAM_832 run on same tapes.

13914 024 08:22 Triggers stop. KFREND fixes problem.

13914 028 08:28 Notice strange auxiliary peak in NM2SEED at
very low end of vertical range: notify MCR.
Possibly this is a hot SEED wire.

13914 037 08:38 Stuck DYC BUSY 4-23 (DC2, SA2)

13914 042 08:46 DYC BUSY unstuck after numerous KFRENDs.

13914 150 11:12 Tapes full: end run.

13915 001 11:23 New BEAM_832 run starts.

13915 177 15:30 end of shift.

========================================================================
Follow up Notes on Previous Shift

We followed the orders of not flipping the analysis magnet field and
not doing a muon run.

========================================================================
Items to Follow up on Next Shift

Keep an eye on the E/P (csigain) of CsI 3114.
A long laser scan is needed if there is downtime.